OpenStudy (anonymous):

I am to solve the exponential equation of 5^(x+3)=242 in natural logarithms. I have so far for the solution set x+3= ln 242/ln 5 I need help better understanding how to do this and on how to use a calculator to give the decimal approximation for the set. Which I have as: .41046876288 How does this look?

OpenStudy (neer2890):

do you want to know how to set the calculator upto some decimal points...?

OpenStudy (anonymous):

I think I have figured that out and now I am trying to see if my solutions are correct. :-)

OpenStudy (anonymous):

Meaning if it is correct I figured it out. :-D

OpenStudy (neer2890):

i don't understand what you want to know...?

OpenStudy (anonymous):

If I have the right logarithm solution set and the right decimal approximation of x in the equation.

OpenStudy (neer2890):

x+3=3.41046876288 x=.41046876288 yeah, definitely,it's correct.

OpenStudy (anonymous):

Thanks. :-) Do you know how to properly express the solution set? Because I am unsure. I think it is to be \[x+3=\frac{ \ln 242 }{ \ln 5 }\]

OpenStudy (neer2890):

this property you can use to find the value of log in any base. \[\ln _{5}242=\frac{ \ln 242 }{ \ln 5 }\]

OpenStudy (anonymous):

Thanks. :-D

OpenStudy (neer2890):

any time... @ShortStaticBurst ...:)